# Two prover perfect zero knowledge for MIP*

### v1

@@ 8. Perfect zero knowledge
Let ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ be a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol for a language $\mathcal{L}$ with completeness $1$ and soundness $1-f(x)$, such that each context of $B_{x}$ has constant size, and $\pi_{x}$ is maximized on the diagonal. Then there is a $\operatorname{PZK}$-$\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})\},\widetilde{S},\widetilde{C})$ for $\mathcal{L}$ with completeness $1$ and soundness $1-f(x)/\operatorname{poly}(m_{x})$, where $m_{x}$ is the number of contexts in $B_{x}$. If $\pi_{x}$ is uniform, then $\pi_{x}^{\prime}$ is also uniform.
@@
Let $B_{x}^{\prime}=\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$, and let $\pi_{x}^{\prime}$ be the subdivision of $\pi_{x}$ corresponding to the subdivision of $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ into $\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$. If $\pi_{x}$ is uniform, then $\pi_{x}^{\prime}$ is also uniform. For completeness, if there’s a perfect tracial state on $\mathcal{A}(B_{x})$, then there is a perfect tracial state on $\mathcal{A}(\operatorname{Obl}_{5}(B_{x}))$ by Lemma 8.4, and consequently a perfect tracial state on $\mathcal{A}(\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x})))$ by Lemma 8.9. By 6.4, there is a perfect tracial state on $\mathcal{A}(B_{x}^{\prime})$. Hence if $x\in\mathcal{L}$, then $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ has a perfect strategy.
@@
Because $B_{x}$ has contexts of constant size, $\operatorname{Obl}_{5}(B_{x})$ and hence $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ also has contexts of constant size. As a result, the number and size of the clauses in the constraints of $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ are also constant. We conclude that the parameters $C$, $M$, and $K$ in Theorem 6.5 when going from $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ to $\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$ are all constant. Since $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ has $m_{x}$ contexts, if $\tau$ is a tracial state on $\mathcal{A}(B_{x}^{\prime})$, then there is a tracial state $\tau_{0}$ on $\mathcal{A}(\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x})))$ with $\operatorname{def}(\tau_{0})\leq\operatorname{poly}(m_{x})\operatorname{def}(\tau)$. Since there is a classical homomorphism $\mathcal{A}(B_{x})\to\mathcal{A}(\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x})))$ by Lemmas 8.4 and 8.9, we conclude that there is a tracial state $\tau_{1}$ on $\mathcal{A}(B_{x})$ with $\operatorname{def}(\tau_{1})\leq\operatorname{poly}(m_{x})\operatorname{def}(\tau)$. Hence if $x\not\in\mathcal{L}$, then there is no synchronous strategy $p$ for $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ with $\omega_{q}(\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime}),p)\geq 1-f(n)/\operatorname{poly}(m_{x})$.
Because all the contraints in $B_{x}$ have constant size, it is not hard to see that the Turing machines $S$ and $C$ can be turned into Turing machines $\widetilde{S}$ and $\widetilde{C}$ such that ($(\{\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})\},\widetilde{S},\widetilde{C})$ is a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol for $\mathcal{L}$.
To prove that this protocol is perfect zero knowledge, we need to find a polynomial time simulator $M_{x}$ which samples a correlation $p_{x}(a,b|i,j)$ that is perfect for the tableau game. Furthermore, $p_{x}(a,b|i,j)$ must be a quantum correlation if and only if $x$ is an accept instance of the tableau game.
The tableau game involves the verifier requesting from each prover exactly one of the constraints (1)-(4) from Definition 8.6, and checking their answers for consistency. The simulator $M_{x}$ can efficiently sample any element $z$ from the clauses of $\operatorname{Obl}_{5}(B_{x})$ of the first row of the corresponding tableau by uniformly sampling from {±$\{\pm 1\}$. Elements of the tableau $T_{i}(p,q)$ and randomizers $r_{i}(p,q)$ can be sampled efficiently by uniformly sampling from $S_{5}$. In this way, $M_{x}$ may efficiently simulate answers to (1) and (2) by sampling the elements on the right side of the equation, and computing the element on the left side. Answers to (3) are simulated by sampling $d-1$ elements of $S_{5}$, where $d$ is the constant depth of the permutation branching program used to construct the tableau, and computing the correct $d^{th}$ entry such that the product of the $d$ elements is equal to $\sigma$, the output of the permutation branching program. Lastly, $M_{x}$ can simulate ansers to (4) by sampling elements of the first row of the tableau uniformly as above (matching any pair that are labeled by the same oblivious variable), and sampling other elements uniformly from $S_{5}$. Thus, simulating the response of an individual player Alice is trivial. The responses from Bob need only be consistent with those of Alice on the overlap, with the remainder of the answer sampled as above. This defines our simulatable correlation $p_{x}(a,b|i,j)$ and our simulator $M_{x}$. It is clear that the correlation $p_{x}(a,b|i,j)$ sampled by $M_{x}$ is perfect for the tableau game. All that remains is to show that $x$ is an accept instance if and only if $p_{x}(a,b|i,j)$ is a quantum correlation.
Suppose that $p_{x}(a,b|i,j)$ is a quantum correlation. Then $x$ is an accept instance, as there is a quantum correlation that allows the players to play the $x$ instance of the tableau game perfectly.
Suppose that $w\in\mathcal{L}$ is an accept instance of the tableau game. Then there is some quantum strategy for the $w$ tableau game such that the players always win. By the gapped soundness of the reduction from 3SAT, this implies that the underlying 3SAT instance has a perfect quantum strategy with observables $\sigma_{i}(x)$ for $x\in V_{i}$. Alice and Bob may now choose any set of oblivious observables $\sigma^{\prime}_{i}(x(1)),\dots,\sigma^{\prime}_{i}(x(5))$ such that the exclusive disjunction of these is $\sigma_{i}(x)$, that is $\sigma^{\prime}_{i}(x(1))\cdots\sigma^{\prime}_{i}(x(5))=\sigma_{i}(x)$. So choose $\sigma^{\prime}_{i}(x(j))$ to be observables that are ±${\pm 1}$ with equal probability for $1\leq j\leq 4$ and let $\sigma^{\prime}_{i}(x(5))=\sigma_{i}(x)\sigma^{\prime}_{i}(x(1))\cdots\sigma^{\prime}_{i}(x(4)),$ and note that The values of any four of the $\sigma_{i}^{\prime}(x(j))$ are efficiently sampleable. To play the tableau game, when Alice and Bob receive their questions $x$ and $y$ respectively, they use auxiliary observables to generate shared uniformly distributed randomizers $r_{i}(p,q)$ and construct the tableaux corresponding to the clauses of $x$ and $y$ according to relations (1) to (5) in Definition 8.6. The value for each element $T_{i}(1,q)$ of row one of the tableau is equally likely to be either element of $\{\pi_{i,1}^{p},\pi_{i,-1}^{p}\}$. Note that the simulator only ever has to sample at most four elements of the first row of a tableau, and only the correlation of five or more of these variables depends on the perfect strategy of $B_{w}$. Each randomizer is an independently uniformly sampled element of $S_{5}$ and thus any element of the second and third rows of the tableau is equally likely to be any element of $S_{5}$. Therefore the correlation generated this way is $p_{w}(a,b|i,j)$. ∎
@@
There is a perfect zero knowledge $\operatorname{BCS}$-$\operatorname{MIP}^{*}(2,1,1,1-1/\operatorname{poly}(n))$ protocol for the halting problem in which the verifier selects questions according to the uniform distribution, the questions have length $\operatorname{polylog}(n)$, and the answers have constant length.
@@
By Theorem 2.2, there is a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,V)$ for the halting problem with constant soundness $s<1$, in which $B_{x}$ has a constant number of contexts and contexts of size $\operatorname{polylog}(|x|)$, and $\pi_{x}$ is the uniform distribution on pairs of contexts. By 5.8, ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ can be turned into a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B^{\prime}_{x},\pi_{x})\},S,C)$ where $B_{x}^{\prime}=(X_{x}^{\prime},\{(W_{i}^{x},D_{i}^{x})\})$, $D_{i}$ is a 3SAT instance with number of clauses polynomial in $|x|$, and $|W^{x}_{i}|$ is polynomial in $|x|$. Then by subdividing the $B_{x}^{\prime}$ into a 3SAT we obtain a 3SAT protocol ($(\{\mathcal{G}(B^{3SAT}_{x},\pi^{3SAT}_{x})\},S,C)$ with number of clauses polynomial in $|x|$, and $\pi^{3SAT}_{x}$ is uniform. The theorem follows from 8.10. ∎
@@
Let ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ be the $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol from Theorem 8.11, so in particular $B_{x}$ has $m_{x}$ contexts, where $m_{x}=\operatorname{poly}(|x|)$, and $\pi_{x}$ is the uniform distribution on $[m_{x}]\times[m_{x}]$. Since the uniform distribution is $1/2m_{x}$-diagonally dominant, Theorem 2.1 implies that ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ has soundness $1-1/\operatorname{poly}(n)$ when considered as a $\operatorname{MIP}^{*}$ protocol. The result follows from 7.1 using a polynomial amount of parallel repetition. ∎
We also have

### v2

@@ 9. Perfect zero knowledge
Let ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ be a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol for a language $\mathcal{L}$ with completeness $1$ and soundness $1-f(x)$, such that each context of $B_{x}$ has constant size, and $\pi_{x}$ is maximized on the diagonal. Then there is a $\operatorname{PZK}$-$\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})\},\widetilde{S},\widetilde{C})$ for $\mathcal{L}$ with completeness $1$ and soundness $1-f(x)/\operatorname{poly}(m_{x})$, where $m_{x}$ is the number of contexts in $B_{x}$. If $\pi_{x}$ is uniform, then $\pi_{x}^{\prime}$ is also uniform, and if $\mathcal{G}(B_{x},\pi_{x})$ has a perfect finite-dimensional tracial state, then so does $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$.
@@
Let $B_{x}^{\prime}=\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$, and let $\pi_{x}^{\prime}$ be the subdivision of $\pi_{x}$ corresponding to the subdivision of $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ into $\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$. If $\pi_{x}$ is uniform, then $\pi_{x}^{\prime}$ is also uniform. Let $p_{x}$ be the correlation for $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ defined in 9.12. Because $B_{x}$ has contexts of constant size, $\operatorname{Obl}_{5}(B_{x})$ and $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ also have contexts of constant size. As a result, the number of clauses in the constraints of $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ is constant, as is the size of each clause (where by clause we mean the constraints of type (1)-(4) in Definition 9.6). Hence the Turing machines $S$ and $C$ can be turned into Turing machines $\widetilde{S}$ and $\widetilde{C}$ such that ($(\{\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})\},\widetilde{S},\widetilde{C})$ is a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol. Similarly, since all the constraints of $\operatorname{Tab}(\operatorname{Obl}_{5}(B))$ have constant size, there is a Turing machine which, given questions and answers $i,j,\phi,\psi$ for $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$, can produce $p_{x}(\phi,\psi|i,j)$ in polynomial time in $i$, $j$, and $x$. Since the number of answers for any question is constant, the correlation $p_{x}$ can be simulated in polynomial time in $x$.
@@
If $x\in\mathcal{L}$, then $B_{x}$ has a perfect strategy in $C_{qa}$, so $p_{x}\in C_{qa}$, and hence $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ has a perfect strategy in $C_{qa}$. Similarly, if $B_{x}$ has a perfect strategy in $C_{q}$, then $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ has a perfect strategy in $C_{q}$ as well. Conversely, suppose that $\tau$ is a tracial state on $\mathcal{A}(B_{x}^{\prime})$. Since the size of contexts and number of clauses in each constraint of $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ are constant, the parameters $C$, $M$, and $K$ in Theorem 7.5 when going from $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ to $\operatorname{Tab}_{sub}(\operatorname{Obl}_{5}(B_{x}))$ are all constant. Since $\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x}))$ has $m_{x}$ contexts, Theorem 7.5 implies that there is a tracial state $\tau_{0}$ on $\mathcal{A}(\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x})))$ with $\operatorname{def}(\tau_{0})\leq\operatorname{poly}(m_{x})\operatorname{def}(\tau)$. Since there is a classical homomorphism $\mathcal{A}(B_{x})\to\mathcal{A}(\operatorname{Tab}(\operatorname{Obl}_{5}(B_{x})))$ by Lemmas 9.4 and 9.8, we conclude that there is a tracial state $\tau_{1}$ on $\mathcal{A}(B_{x})$ with $\operatorname{def}(\tau_{1})\leq\operatorname{poly}(m_{x})\operatorname{def}(\tau)$. Hence if $x\not\in\mathcal{L}$, then there is no synchronous strategy $p$ for $\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})$ with $\omega_{q}(\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime}),p)\geq 1-f(n)/\operatorname{poly}(m_{x})$. Hence ($(\{\mathcal{G}(B_{x}^{\prime},\pi_{x}^{\prime})\},\widetilde{S},\widetilde{C})$ is a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol for $\mathcal{L}$ with soundness $1-f(x)/\operatorname{poly}(m_{x})$. ∎
@@
There is a perfect zero knowledge $\operatorname{BCS}$-$\operatorname{MIP}^{*}(2,1,1,1-1/\operatorname{poly}(n))$ protocol for the halting problem in which the verifier selects questions according to the uniform distribution, the questions have length $\operatorname{polylog}(n)$, and the answers have constant length. Furthermore, if a game in the protocol has a perfect strategy, then it has a perfect synchronous quantum strategy.
@@
By Theorem 3.2, there is a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,V)$ for the halting problem with constant soundness $s<1$, in which $B_{x}$ has a constant number of contexts and contexts of size $\operatorname{polylog}(|x|)$, and $\pi_{x}$ is the uniform distribution on pairs of contexts. Furthermore, if $\mathcal{G}(B_{x},\pi_{x})$ has a perfect strategy, then it has a perfect synchronous quantum strategy. By 6.8, ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ can be turned into a $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol ($(\{\mathcal{G}(B^{\prime}_{x},\pi_{x})\},S,C)$ where $B_{x}^{\prime}=(X_{x}^{\prime},\{(W_{i}^{x},D_{i}^{x})\})$, $D_{i}$ is a 3SAT instance with number of clauses polynomial in $|x|$, and $|W^{x}_{i}|$ is polynomial in $|x|$. Then by subdividing the $B_{x}^{\prime}$ into a 3SAT we obtain a 3SAT protocol ($(\{\mathcal{G}(B^{3SAT}_{x},\pi^{3SAT}_{x})\},S,C)$ with number of clauses polynomial in $|x|$, and $\pi^{3SAT}_{x}$ is uniform. There is a $1$-homomorphism $\mathcal{A}(B^{3SAT}_{x},\mu_{\pi^{3SAT}_{x}})\to\mathcal{A}(B_{x},\mu_{\pi_{x}})$, so if $\mathcal{G}(B_{x},\pi_{x})$ has a perfect synchronous quantum strategy, so does $\mathcal{G}(B^{3SAT}_{x},\pi^{3SAT}_{x})$. The theorem follows from 9.14. ∎
@@
By the discussion after Theorem 3.2, it is enough to show that there is a two-prover one-round perfect zero knowledge $\operatorname{MIP}^{*}$ protocol for the halting problem with completeness $1$, soundness $1/2$, and uniform probability distribution. Let ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ be the $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol from Theorem 9.15, so in particular $B_{x}$ has $m_{x}$ contexts, where $m_{x}=\operatorname{poly}(|x|)$, and $\pi_{x}$ is the uniform distribution on $[m_{x}]\times[m_{x}]$. Since the uniform distribution is $1/2m_{x}$-diagonally dominant, Theorem 3.1 implies that ($(\{\mathcal{G}(B_{x},\pi_{x})\},S,C)$ has soundness $1-1/\operatorname{poly}(n)$ when considered as a $\operatorname{MIP}^{*}$ protocol. The result follows from 8.1 using a polynomial amount of parallel repetition. ∎
As mentioned in the introduction, the proof of Theorem 1.1 implies that the halting problem is many-one reducible to membership in $C_{q}$. In fact, there is a reduction such that if the Turing machine does not halt, then the corresponding correlation is bounded away from the closure $C_{qa}$ of $C_{q}$:
Corollary 9.16.
There is a polynomial-time computable function $p$ from Turing machines to synchronous correlations such that if $M$ halts then $p(M)\in C_{q}^{s}$, and if $M$ does not halt then there is a linear functional $f$ on the space of correlations such that $f(p(M))=1$ and $f(p^{\prime})\leq 1/2$ for all $p^{\prime}\in C_{qa}$.
Proof.
Let ($(\{\mathcal{G}(B_{M},\pi_{M})\},S,C)$ be the $\operatorname{BCS}$-$\operatorname{MIP}^{*}$ protocol for the halting problem with completeness one and soundness $1/2$ constructed in the proof of Theorem 1.1, where the index $M$ runs through Turing machines. Let $p(M)$ be the correlation for $\mathcal{G}(B_{M},\pi_{M})$ as in Definition 9.1. That $p(M)$ is in $C_{q}$ follows from Theorem 9.15, and the fact that if $p\in C_{q}$, then $p^{\otimes n}\in C_{q}$. The corollary then follows with the linear functional $f$ defined by $f(p^{\prime})=\omega(\mathcal{G}(B_{M},\pi_{M}),p^{\prime})$. ∎
Note that the number of inputs and outputs for the correlation $p(M)$ depends on the size of the Turing machine $M$.
Finally, we also have: